# 11Be; halo structure; ab initio calculation; nuclear lattice effective field theory

### 07 Mar 2026 v1

@@ 3 Results and discussion
In Ref. Shen:2024qzi, the $A$-nucleons are grouped into 2 clusters composed of 2 protons and 2 neutrons that are the closest to each other according to their spatial coordinates $\mathbf{r}_{i}$ sampled by the pinhole algorithm. The extra neutrons (2 in 10Be and 3 in 11Be) are considered to be the valence neutrons. By randomly rotating one of the clusters along the $z$-axis, the intrinsic 3D density distribution $\rho(\mathbf{r})$ is obtained. The results of 10Be and 11Be ground states are shown in Fig. 11(a) and (b), respectively. Their difference (11Be - 10Be) is shown in Fig. 1(c).
@@ 4 Conclusions and Perspectives
Conceptualization, S.S. and U.G.M.; methodology, all authors; software, all authors; validation, all authors; formal analysis, S.S.; investigation, S.S.; resources, S.S. and U.G.M.; data curation, S.S. and U.G.M.; writing—original draft preparation, S.S.; writing—review and editing, all authors; visualization, S.S.; supervision, U.G.M.; project administration, U.G.M.; funding acquisition, S.S. and U.G.M. All authors have read and agreed to the published version of the manuscript.
@@
This work was funded by “the Fundamental Research Funds for the Central Universities”, National Natural Science Foundation of China under Grant No. U2541242 and 12435007. The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of UGM was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). SS gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University. SS and UGM gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). DL acknowledges support from the U.S. Department of Energy grants DE-SC0026198, DE-SC0013365, and DE-SC0023175.

### 11 Mar 2026 v2

@@ 3 Results and discussion
In Ref. Shen:2024qzi, the AA-nucleons are grouped into 2 clusters composed of 2 protons and 2 neutrons that are the closest to each other according to their spatial coordinates $\mathbf{r}_{i}$ sampled by the pinhole algorithm. The extra neutrons (2 in 10Be and 3 in 11Be) are considered to be the valence neutrons. By randomly rotating one of the clusters along the $z$-axis, the intrinsic 3D density distribution $\rho(\mathbf{r})$ is obtained. The results of 10Be and 11Be ground states are shown in Fig. 11(a) and (b), respectively. Their difference (11Be - 10Be) is shown in Fig. 1(c).
@@ 4 Conclusions and Perspectives
Conceptualization, $S.$ and U.G.M.; methodology, all authors; software, all authors; validation, all authors; formal analysis, $S.$; investigation, $S.$; resources, $S.$ and U.G.M.; data curation, $S.$ and U.G.M.; writing—original draft preparation, $S.$; writing—review and editing, all authors; visualization, $S.$; supervision, U.G.M.; project administration, U.G.M.; funding acquisition, $S.$ and U.G.M. All authors have read and agreed to the published version of the manuscript.
@@
This work was funded by “the Fundamental Research Funds for the Central Universities”, National Natural Science Foundation of China under Grant No. U2541242 and 12435007. The work of SE is supported in part by the Scientific and Technological Research Council of Turkey (TUBITAK Project No. 123F464). The work of UGM was supported in by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (ERC AdG EXOTIC, Grant Agreement No. 101018170), by the CAS President’s International Fellowship Initiative (PIFI) (Grant No. 2025PD0022). $S$ gratefully acknowledge the Computational resources provided by the HPC platform of Beihang University. $S$ and UGM gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). DL acknowledges support from the U.S. Department of Energy grants DE-SC0026198, DE-SC0013365, and DE-SC0023175.